If a business borrows $24,000 and repays $34,500 in 5 years, what is the simple interest rate?

Answer by rfer(16322)

34500-24000=10500
I=Prt
r=I/Pt
r=10500/24000*5
r=10500/120000
r=0.0875=8.75%